if your distance from the foot of the tower is 20m and the angle of elevation is 40 degrees, find the height of the tower?
Answer by jim_thompson5910(35256)

tan(angle) = opp/adj

tan(40) = x/20

20*tan(40) = x

x = 20*tan(40)

x = 16.7819926

So the tower is roughly 16.7819926 meters tall.